# Diffusion-HMC: Parameter Inference with Diffusion Model driven Hamiltonian Monte Carlo

### v1

@@ Abstract
Diffusion generative models have excelled at diverse image generation and reconstruction tasks across fields. A less explored avenue is their application to discriminative tasks involving regression or classification problems. The cornerstone of modern cosmology is the ability to generate predictions for observed astrophysical fields from theory and constrain physical models from observations using these predictions. This work uses a single diffusion generative model to address these interlinked objectives – as a surrogate model or emulator for cold dark matter density fields conditional on input cosmological parameters, and as a parameter inference model that solves the inverse problem of constraining the cosmological parameters of an input field. The model is able to emulate fields with summary statistics consistent with those of the simulated target distribution.
We then leverage the approximate likelihood of the diffusion generative model to derive tight constraints on cosmology by using the Hamiltonian Monte Carlo method to sample the posterior on cosmological parameters for a given test image. Finally, we demonstrate that this parameter inference approach is more robust to the addition of noise than baseline parameter inference networks.
@@ I Introduction
The canonical summary statistic used for parameter inference is the two point correlation function, or the power spectrum $Pk$ at large, linear scales where perturbation theory holds. At smaller scales, however, gravitational collapse induces non-Gaussianity in the fields. This means the information content of cosmological fields is not fully captured by the power spectrum at the large scales alone. For example, [11] derived 27% stronger constraints of $\sigma_{8}$ from the two point correlation function of galaxy clustering data by analyzing non-linear modes at smaller scales, ($k>$0.25 $h\rm{Mpc}^{-1}$). A multitude of other statistics – such as the marked power spectrum, the bispectrum, the wavelet scattering transform, and void probability functions – have also been devised [44, 28, 32, 12] in an effort to capture higher order correlations in non-Gaussian fields. Previous work [14, 23, 28, 45, 38] has addressed the prohibitive cost of simulations by creating emulators or surrogate models that learn to interpolate between predictions of a specific summary statistic between training points using formalisms such as Gaussian processes. More recently, simulation-based inference at the field level has also been used, where a likelihood model parameterized by a neural network is learned on the fields [9, 8].
@@
Generative models are a class of machine learning approaches that enable one to simulate the ability to draw samples from a complicated target probability density and include variational autoencoders, normalizing flows and generative adversarial networks (GANs). Diffusion generative models [41] involve a forward diffusion (noising) process that transform samples from the target distribution to those from the standard normal. In the denoising diffusion probabilistic model (DDPM)[16], the noising process consists of a variance schedule ${\beta_{t}}$ over a fixed number of time steps, $T$, that determines the incremental noise added to the image. Since the diffusion process can be formulated as a stochastic differential equation (SDE) the DDPM variance schedule corresponds to the discretization of this SDE. In the generative direction, a neural network or score model is used to parameterize the reverse transformation.
@@
Diffusion models are alternatively referred to as a score-based generative model since parameterizing the reverse diffusion process is equivalent to to learning the ‘score’ or $\vec{\nabla}\log p_{t}(x)$ of the data [1, 41]. Since in high dimensions, the target distribution invariably lies on a thin manifold, the incremental addition of the random noise, blurs the distribution and makes the score progressively easier to learn. Diffusion models are the underlying mathematical framework that have given rise to the photorealistic image generation successes of DALL.E and Stable Diffusion [34] and have been shown to mitigate mode collapse, a phenomenon, often encountered with GANs a generative model fails to generate multiple modes in a distribution. In scientific applications, they have used for problems involving protein folding and ligand prediction and medical imaging reconstruction [7, 40] . They have been applied in astrophysics to reconstruction problems involving dust [15, 22], cosmological simulations and initial conditions reconstruction [20, 36, 27, 8] and strong lensing problems [18, 33].
@@
In this work, we apply diffusion generative models to emulating cold dark matter density fields conditional on cosmological parameters and demonstrate that the trained model can also be used to derive tight and robust constraints on cosmological parameters. In Section III, we examine the ability of the model to appropriately capture the statistics of the distribution of fields corresponding to different parameters, and further compare the effect of modulating a single parameter at a time on the statistics of the resulting fields in the true and the generated set. We then quantify the ability of the model to capture the full range of cosmic variance for a single parameter, as a means to assess the extent of mode collapse. In Section  IV, we examine how the diffusion model’s approximate likelihood can be used to solve the inverse problem of constraining the cosmological parameters of a given input field. We then use Hamiltonian Monte Carlo [10, 24, 3] to draw samples from the estimated posterior on the cosmological parameters given an input field, and compare our estimates with a power spectrum baseline. A novel contribution of this work is our use of an HMC to sample a posterior consisting of an approximation to the diffusion model conditional likelihood to solve a downstream inference task. Finally, we demonstrate that the Diffusion-HMC based parameter inference estimates are more robust to the addition of uncorrelated noise relative to the estimates from a discriminative neural network directly trained to estimate parameters.
@@ II Datasets, Architecture and Training
We work with cold Dark Matter density fields at $z=0$ from the IllustrisTNG [25, 30] suite from the CAMELS Multifield Dataset (CMD) [49, 47]. The dataset contains of 1000 simulations for 1000 different cosmologies with 15 two-dimensional fields per simulation. The ‘cosmology’ is parameterized by a parameter vector with 2 cosmological ($\Omega_{m}$ and $\sigma_{8}$) and 4 astrophysical parameters. Each simulation tracked the evolution of $256^{3}$ dark matter particles and $256^{3}$ fluid elements and took around 6000 CPU hours to generate (see [49, 47] for more details). The fields span $25$ Mpc.$h^{-1}$ on each side. We train on 70% of the parameters in the LH set, i.e. 700 parameters or 10,500 fields, and condition the model only on the cosmological parameters $\Omega_{m}$ and $\sigma_{8}$. Example generated dark matter fields are shown in Figure  2.
@@
The score / noise-predictor model is a U-Net [35] similar to that used in [16]. There are 4 downsampling layers, with each layer consisting of 2 ResNet blocks [53], group-normalization [52], and attention [46, 39]. We use circular convolutions in the downsampling layers since the input fields have periodic boundary conditions. Each parameter is normalized to lie between [0, 1] with respect to its range, $\Omega_{m}\in[0.1,0.5],\sigma_{8}\in[0.6,1.0]$. A multilayer perceptron (MLP)transforms the cosmology vector into a space with the same dimension as the time embedding, and each ResNet block additionally has an MLP conditional on cosmology. The variance schedule $\beta_{t}$ is non-linear with smaller steps at smaller $t$ and larger steps for larger values of $t$ (see Figure  8). We train the model to generate the log of the fields, and randomly rotate and flip the image to account for these invariances. During training, for each batch of images $\mathbf{x}_{0}$, a batch of timesteps is sampled uniformly along with a noise pattern $\epsilon\sim\mathcal{N}(0,\mathbf{I})$. The loss function minimized is $||\mathbf{\epsilon}-\mathbf{\epsilon}_{\phi}(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon},t,\theta))||^{2}$ for each set of $\{\mathbf{x}_{0},t,\theta,\epsilon\}$, where $\theta$ is the parameter vector. In practice, we minimize the Huber loss, which behaves as an L1 loss for larger values and L2 for smaller values of the loss. We used the Weights and Biases framework [4] to track experiments. The model has 31.2 million parameters.
@@
The noise prediction loss does not fully capture sample quality and convergence, and we need an alternative metric to assess the quality of the generated samples [43]. We sampled 500 fields (with 50 fields for 10 different validation parameters) for the checkpoints after 200k, 220k, 240k, 260k, 280k, 300k, 320k and 340k iterations. We computed the reduced chi-squared statistic (Equation  10) of the power spectrum of each generated field, $s$, with respect to the reference distribution comprised of the 15 true fields for that parameter. We then compute the mean and standard error of these values across all parameters and sampled fields.
While the diffusion generative model is trained to generate the fields in log space, the power spectra we compute here and in Figures  13 correspond to the overdensity power spectra of the ‘linear’ ($10^{\rm{GeneratedFields}}$) fields. The checkpoint corresponding to the 260kth iteration had the lowest value, corresponding to $2.29\pm 0.49$. To put this number in perspective, we can examine the effect of cosmic variance on this metric using a leave-one-out cross-validation approach, by computing the reduced chi-squared statistic of each sample of a true field, using the 14 other true fields corresponding to the same parameter as the reference distribution. The mean of this value across the 10 parameters is $1.70\pm 0.36$. We use the 260k checkpoint for our analysis. We plot these values, along with the reduced chi-squared statistic of the power spectra of the log fields and the $p$ values of the mean intensity in Appendix .3.1.
@@ III Summary Statistics
In this section we examine the consistency of our generated fields relative to the true fields using three sets of simulations under the IllustrisTNG CAMELS suite. The latinhypercube, LH, suite varies the largest range of cosmological and astrophysical parameters but has a limited number (15) of fields for each parameter, with all parameters varying randomly. The one-parameter, 1P, suite consists of 15 fields with the same seed, and one dimensional variations of each parameter modulated systematically, while the others are fixed to the fiducial value. The Cosmic Variance, CV, set consists of 405 fields for the fiducial parameter value. Sampling a batch of 50 $256\times 256$ fields from our model takes 310 seconds (6.2s/field).
@@
Varying cosmological parameters one at a time (1P): In Figure  2, we generated ‘1P’ sets and examined whether the effect of modulating a single parameter, while keeping the others constant is the same as is observed in the 1P CAMELS suite. We sample 15 fields corresponding to 15 different seeds for each of the parameters. The fields corresponding to the same seed across parameters have the same position and orientation of their seeded structures as visible in Figure  2. For each seed, we then compute the ratio of the power spectrum of a field at a different parameter to the power spectrum of the fiducial parameter value, and compute the average and the percentile-based standard deviation envelopes across all seeds as depicted in the right column of Figure  2. The dashed (solid) line and envelope correspond to the generated (true) ratios. The ratios for the generated ‘1P’ set are in good agreement with the those of the true ‘1P’ set, since the dashed lines are within the solid envelopes.
@@
The standard deviations of the distribution are also consistent although the standard deviations of the generated power spectra are slightly underpredicted relative to the true power spectra at the largest length scales ($k<5$$h\rm{Mpc}^{-1}$). To construct the standard deviation estimator, we use the jackknife-approach to generate a distribution of 450 estimates of the standard deviation for each set, where the $\rm{i^{th}}$ estimate corresponds to the standard deviation computed using all samples excluding that of the $\rm{i^{th}}$ field. We then compute the mean and the standard deviation for these sets in order to capture the mean and the standard deviation on the estimate of the standard deviation in the upper right panel of Figure  3. We use the ratios of the jackknife-estimated means and errors in the lower right panel.
@@
Moreover, the high dimensionality of the initial conditions relative to our training set size of 10,000 fields poses a challenge. Increasing the size of the training dataset may help mitigate this issue.
@@ IV Parameter Inference
$\mathbb{E}_{q}[-\log p_{\phi}(x_{0}|\theta)]\leq\mathbb{E}_{q}[{\color[rgb]{0,0,0}-\log p_{\phi}(x_{0}|x_{1},\theta)}+$
${\color[rgb]{0,0,0}\sum_{t\geq 1}\rm{D_{KL}}[q(x_{t}|x_{t+1},x_{0})||p_{\phi}(x_{t}|x_{t+1},\theta)]}+$
${\color[rgb]{0,0,0}\rm{D_{KL}}[q(x_{T}|x_{0})||p(x_{T})]}]$ (2)
$L_{\rm{VLB}}={\color[rgb]{0,0,0}L_{0}}+{\color[rgb]{0,0,0}L_{1}...L_{T-1}}+{\color[rgb]{0,0,0}L_{T}}$ (3)
@@ IV.2 HMC-based Parameter Inference
To compute the parameter estimates for fields, we draw samples from the posterior on the parameter using the Hamiltonian Monte Carlo (HMC) method. The Hamiltonian Monte Carlo, or Hybrid Monte Carlo method [10, 24, 3] draws samples from a probability distribution $\pi(\theta)$ via the introduction of an auxiliary momentum variable $\mathbf{p}$ and solves the equations of Hamiltonian dynamics in order to update the momentum and position ($\theta$). HMC enables a more efficient exploration of high dimensional probability distributions. In our case, using an HMC also helps us circumvent the problem of having to redefine the extents and granularity of the parameter grid depending on how confident the constraints for a given $T_{\rm{MAX}}$ are. The Hamiltonian governing the dynamics in the HMC chain is:
@@
We used the Hamiltorch [6] package for the HMC. We explain more details about the HMC implementation in Appendix  .1. The prior is chosen to be a flat prior over $\Omega_{m}\in[0.1,0.5]$ and $\sigma_{8}\in[0.6,1.0]$. The initial parameter value is always the fiducial value of $[0.3,0.8]$ and we designate the first 100 samples as burn-in samples that are discarded.
We now examine the effect of truncating terms in Equation  4 using the HMC based parameter inference. Truncating terms allows us to perform inference faster and can allow us to explore the tradeoff between dropping terms for speed and higher precision with more timesteps. In Figure  5b), for a single field, we plot the mean predictions and the $15.9-84.1$ (1 sigma) percentiles computed using 200 samples with the approximate $-\log p_{\phi}(x_{0}|\theta)$ using Equation  4 as a function of $T_{\rm{MAX}}$ on the x axis for the same field. For this field and parameter, using more terms asymptotically removes the bias on $\Omega_{m}$ while increasing the bias on $\sigma_{8}$. However, using the first 20 timesteps only changes the mean prediction for $\Omega_{m}$ and $\sigma_{8}$ by $-0.26\%$ and $-1.27\%$ of the prediction using all 1000 timesteps respectively.
Refer to captionRefer to captionRefer to caption
@@
We now turn our attention to the performance of our parameter inference approach relative to a power spectrum baseline. For subsequent HMC-based parameter inference in this section, we use Equation  4 with $T_{\rm{MAX}}=20$ to approximate the conditional negative log likelihood. Drawing 500 samples with $T_{\rm{MAX}}=20$ takes $\sim 32$ minutes for a single field.
To assess the constraining power of the power spectrum, we use the neural posterior estimation approach and train a normalizing flow to represent the posterior of $p_{\rm{NF}}(\theta|Pk\cdot\rm{Mean})$ using the Lampe [37] package. Since computing the overdensity power spectrum involves dividing by the mean of the field, we further concatenate the mean of the log field as an additional feature, since the prediction for $\Omega_{m}$ for a small box size of $25$ Mpc.$h^{-1}$ is very sensitive to the mean of the fields. For a single field, we compare the posteriors obtained by drawing 10000 samples from the power spectrum based estimator and 400 samples from the diffusion model+HMC in Figure  6. Other details about our implementation are in the Appendix  .2.
@@
Figure 7: Mean and $15.8-84.1$ percentile predictions using the Diffusion-HMC estimates (green) and the mean and sigma predictions obtained from the parameter inference network in [47], denoted by InfNet. The mean absolute percentage bias (of the true prediction) for each parameter is also indicated. The Diffusion-HMC constraints are more robust to the perturbations to the input image.
IV.2.1 Robustness

### v2

# Diffusion-HMC: Parameter Inference with Diffusion-model-driven Hamiltonian Monte Carlo

@@ Abstract
Diffusion generative models have excelled at diverse image generation and reconstruction tasks across fields. A less explored avenue is their application to discriminative tasks involving regression or classification problems. The cornerstone of modern cosmology is the ability to generate predictions for observed astrophysical fields from theory and constrain physical models from observations using these predictions. This work uses a single diffusion generative model to address these interlinked objectives – as a surrogate model or emulator for cold dark matter density fields conditional on input cosmological parameters, and as a parameter inference model that solves the inverse problem of constraining the cosmological parameters of an input field. The model is able to emulate fields with summary statistics consistent with those of the simulated target distribution. We then leverage the approximate likelihood of the diffusion generative model to derive tight constraints on cosmology by using the Hamiltonian Monte Carlo method to sample the posterior on cosmological parameters for a given test image. Finally, we demonstrate that this parameter inference approach is more robust to small perturbations of noise to the field than baseline parameter inference networks.
Astrostatistics (1882), Cosmological parameters (339), Astronomical simulations (1857), Cosmology (343), Dark matter (353), Neural networks (1933)
@@ I Introduction
The canonical summary statistic used for parameter inference is the two point correlation function, or the power spectrum $Pk$ at large, linear scales where perturbation theory holds. At smaller scales, however, gravitational collapse induces non-Gaussianity in the fields. This means the information content of cosmological fields is not fully captured by the power spectrum at the large scales alone. For example, recent work [12, 30] derived much stronger constraints on $\sigma_{8}$ by going beyond the two point correlation function at linear scales and analyzing non-linear modes at smaller scales ($k>$0.25 $h\rm{Mpc}^{-1}$) or by extracting information at the field-level. A multitude of other statistics – such as the marked power spectrum, the bispectrum, the wavelet scattering transform, and void probability functions – have also been devised [51, 33, 37, 13] in an effort to capture higher order correlations in non-Gaussian fields. Previous work [15, 27, 33, 52, 44] has addressed the prohibitive cost of simulations by creating emulators or surrogate models that learn to interpolate between predictions of a specific summary statistic between training points using formalisms such as Gaussian processes. More recently, simulation-based inference at the field level has also been used, where a likelihood model parameterized by a neural network is learned on the fields [10, 9].
@@
Generative models are a class of machine learning approaches that enable one to simulate the ability to draw samples from a complicated target probability density and include variational autoencoders, normalizing flows and generative adversarial networks (GANs). Diffusion generative models [46, 48] involve a forward diffusion (noising) process that transform samples from the target distribution to those from the standard normal. In the denoising diffusion probabilistic model (DDPM)[17], the noising process consists of a variance schedule ${\beta_{t}}$ over a fixed number of time steps, $T$, that determines the incremental noise added to the image. Since the diffusion process can be formulated as a stochastic differential equation (SDE) the DDPM variance schedule corresponds to the discretization of this SDE. In the generative direction, a neural network or score model is used to parameterize the reverse transformation.
@@
Diffusion models are alternatively referred to as score-based generative models since parameterizing the reverse diffusion process is equivalent to learning the ‘score’ or $\vec{\nabla}\log p_{t}(x)$ of the data [1, 48]. Since in high dimensions, the target distribution invariably lies on a thin manifold, the incremental addition of the random noise, blurs the distribution and makes the score progressively easier to learn. Diffusion models are the underlying mathematical framework that have given rise to the photorealistic image generation successes of DALL.E and Stable Diffusion [39] and have been shown to mitigate mode collapse, a phenomenon, often encountered with GANs a generative model fails to generate multiple modes in a distribution. In scientific applications, they have been used for problems involving protein folding and ligand prediction and medical imaging reconstruction [8, 47] . They have been applied in astrophysics to reconstruction problems involving dust [16, 26], cosmological simulations and initial conditions reconstruction [23, 41, 32, 9, 25] and strong lensing problems [19, 38].
@@
In this work, we apply diffusion generative models to emulating cold dark matter density fields conditional on cosmological parameters and demonstrate that the trained model can also be used to derive tight and robust constraints on cosmological parameters. In Section III, we examine the ability of the model to appropriately capture the statistics of the distribution of fields corresponding to different parameters, and further compare the effect of modulating a single parameter at a time on the statistics of the resulting fields in the true and the generated set. We then quantify the ability of the model to capture the full range of cosmic variance for a single parameter, as a means to assess the extent of mode collapse. In Section  IV, we examine how the diffusion model’s approximate likelihood can be used to solve the inverse problem of constraining the cosmological parameters of a given input field. We then use the Hamiltonian Monte Carlo [11, 28, 3] method to draw samples from the estimated posterior on the cosmological parameters given an input field, and compare our estimates with a power spectrum baseline. A novel contribution of this work is our use of an HMC to sample a posterior consisting of an approximation to the diffusion model conditional likelihood to solve a downstream inference task. Finally, we demonstrate that the Diffusion-HMC based parameter inference estimates are more robust to perturbations composed of uncorrelated noise relative to the estimates from a discriminative neural network directly trained to estimate parameters.
@@ II Datasets, Architecture and Training
We work with cold Dark Matter density fields at $z=0$ from the IllustrisTNG [29, 35] suite from the CAMELS Multifield Dataset (CMD) [56, 54]. The diffusion model is trained to generate the minmax transform applied to the log (base 10) of these dark matter fields. The minmax transform is pegged to the minimum and maximum of the log of the entire dataset, [9.42, 15.44]. The dataset contains 1000 simulations for 1000 different cosmologies with 15 two-dimensional fields per simulation. The ‘cosmology’ is parameterized by a parameter vector with 2 cosmological ($\Omega_{m}$ and $\sigma_{8}$) and 4 astrophysical parameters. Each simulation tracked the evolution of $256^{3}$ dark matter particles and $256^{3}$ fluid elements and took around 6000 CPU hours to generate (see [56, 54] for more details). The fields span $25$ Mpc.$h^{-1}$ on each side. We train on 70% of the parameters in the LH set, i.e. 700 parameters or 10,500 fields, and condition the model only on the cosmological parameters $\Omega_{m}$ and $\sigma_{8}$. Since we use dark matter density fields in this study, we do not expect them to constrain or contain much information about the astrophysical parameters. Example generated dark matter fields are shown in Figure  2.
@@
The score / noise-predictor model is a U-Net [40] similar to that used in [17]. There are 4 downsampling layers, with each layer consisting of 2 ResNet blocks [60], group-normalization [59], and attention [53, 45]. We use circular convolutions in the downsampling layers since the input fields have periodic boundary conditions. Each parameter is normalized to lie between [0, 1] with respect to its range, $\Omega_{m}\in[0.1,0.5],\sigma_{8}\in[0.6,1.0]$. A multilayer perceptron (MLP) transforms the cosmology vector into a space with the same dimension as the time embedding, and each ResNet block additionally has an MLP conditional on cosmology. The variance schedule $\beta_{t}$ is non-linear with smaller steps at smaller $t$ and larger steps for larger values of $t$ (see Figure  8). We train the model to generate the log of the fields, and randomly rotate and flip the image to account for these invariances. During training, for each batch of images $\mathbf{x}_{0}$, a batch of timesteps is sampled uniformly along with a noise pattern $\epsilon\sim\mathcal{N}(0,\mathbf{I})$. The loss function minimized is $||\mathbf{\epsilon}-\mathbf{\epsilon}_{\phi}(\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\mathbf{\epsilon},t,\theta))||^{2}$ for each set of $\{\mathbf{x}_{0},t,\theta,\epsilon\}$, where $\theta$ is the parameter vector. Our implementation minimizes the Huber loss, which behaves as an L1 loss for values of the loss greater than 1 and a Mean Squared Error (MSE) loss otherwise. From the training curves, the loss during training is less than 1 throughout, so the loss being minimized is in effect the MSE loss. We used the Weights and Biases framework [4] to track experiments. The model has 31.2 million parameters.
@@
The noise prediction loss does not fully capture sample quality and convergence, and we need an alternative metric to assess the quality of the generated samples [50]. We sampled 500 fields (with 50 fields for 10 different validation parameters) for the checkpoints after 200k, 220k, 240k, 260k, 280k, 300k, 320k and 340k iterations. Sampling a batch of 50 $256\times 256$ fields from our model takes 310 seconds (6.2s/field). We computed the reduced chi-squared statistic (Equation  10) of the power spectrum of each generated field, $s$, with respect to the reference distribution comprised of the 15 true fields for that parameter. We then compute the mean and standard error of these values across all parameters and sampled fields.
While the diffusion generative model is trained to generate the fields in log space, the power spectra we compute here and in Figures  13 correspond to the overdensity power spectra of the ‘linear’ ($10^{\rm{GeneratedFields}}$) fields. The checkpoint corresponding to the 260kth iteration had the lowest value for the chi-squared statistic of the power spectra of the linear generated fields relative to the distribution of true fields, corresponding to $2.29\pm 0.49$ (Figure 12). To put this number in perspective, we can examine the effect of cosmic variance on this metric using a leave-one-out cross-validation approach, by computing the reduced chi-squared statistic of each sample of a true field, using the 14 other true fields corresponding to the same parameter as the reference distribution. The mean of this value across the 10 parameters is $1.70\pm 0.36$. We use the 260k checkpoint for our analysis. We plot these values, along with the reduced chi-squared statistic of the power spectra of the log fields and the $p$ values of the mean intensity in Appendix B.1.
@@ III Summary Statistics
In this section we examine the consistency of our generated fields relative to the true fields using three sets of simulations under the IllustrisTNG CAMELS suite. The Latin Hypercube, LH, suite varies the largest range of cosmological and astrophysical parameters but has a limited number (15) of fields for each parameter, with all parameters varying randomly. The one-parameter, 1P, suite consists of 15 fields with the same seed, and one dimensional variations of each parameter modulated systematically, while the others are fixed to the fiducial value. The Cosmic Variance, CV, set consists of 405 fields for the fiducial parameter value.
@@
Varying cosmological parameters one at a time (1P): In Figure  2, we generated ‘1P’ sets and examined whether the effect of modulating a single parameter, while keeping the others constant is the same as is observed in the 1P CAMELS suite. We sample 15 fields corresponding to 15 different seeds for each of the parameters. The fields corresponding to the same seed across parameters have the same position and orientation of their seeded structures as visible in Figure  2. For each seed, we then compute the ratio of the power spectrum of a field at a different parameter to the power spectrum of the fiducial parameter value, and compute the average and the percentile-based standard deviation envelopes across all seeds as depicted in the right column of Figure  2. The dashed (solid) line and envelope correspond to the generated (true) ratios. The ratios for the generated ‘1P’ set are in good agreement with those of the true ‘1P’ set, since the dashed lines are within the solid envelopes.
@@
The standard deviations of the distribution are also consistent although the standard deviations of the generated power spectra are slightly underpredicted relative to the true power spectra at the largest length scales ($k<5$$h\rm{Mpc}^{-1}$). To construct the standard deviation estimator, we use the jackknife-approach to generate a distribution of 405 estimates of the standard deviation for each set, where the $\rm{i^{th}}$ estimate corresponds to the standard deviation computed using all samples excluding that of the $\rm{i^{th}}$ field. We then compute the mean and the standard deviation for these sets in order to capture the mean and the standard deviation on the estimate of the standard deviation in the upper right panel of Figure  3. We use the ratios of the jackknife-estimated means and errors in the lower right panel.
@@ IV Parameter Inference
$\mathbb{E}_{q}[-\log p_{\phi}(x_{0}|\theta)]\leq\mathbb{E}_{q}[{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}-\log p_{\phi}(x_{0}|x_{1},\theta)}+$
${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\sum_{t\geq 1}\rm{D_{KL}}[q(x_{t}|x_{t+1},x_{0})||p_{\phi}(x_{t}|x_{t+1},\theta)]}+$
${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\rm{D_{KL}}[q(x_{T}|x_{0})||p(x_{T})]}]\simeq L_{\rm{VLB}}$ (2)
$L_{\rm{VLB}}={\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}L_{0}}+{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}L_{1}...L_{T-1}}+{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}L_{T}}$ (3)
@@ IV.2 HMC-based Parameter Inference
To compute the parameter estimates for fields, we draw samples from the posterior on the parameter using the Hamiltonian Monte Carlo (HMC) method. The Hamiltonian Monte Carlo, or Hybrid Monte Carlo method [11, 28, 3] is a Markov Chain Monte Carlo (MCMC) approach that draws samples from a probability distribution $\pi(\theta)$ via the introduction of an auxiliary momentum variable $\mathbf{p}$ and solves the equations of Hamiltonian dynamics in order to update the momentum and position ($\theta$). HMC enables a more efficient exploration of high dimensional probability distributions. In our case, using an HMC also helps us circumvent the problem of having to redefine the extents and granularity of the parameter grid depending on how confident the constraints for a given $T_{\rm{MAX}}$ are. The Hamiltonian governing the dynamics in the HMC chain is:
@@
We used the Hamiltorch [7] package for the HMC. We explain more details about the HMC implementation in Appendix  8. The prior is chosen to be a flat prior over $\Omega_{m}\in[0.1,0.5]$ and $\sigma_{8}\in[0.6,1.0]$. The initial parameter value is always the fiducial value of $[0.3,0.8]$ and we designate the first 100 samples as burn-in samples that are discarded. We reduce the averages over random noise in each $L_{t}$ term in Equation 2 to a single stochastic estimate of $L_{t}$ in Equation 4.
We now examine the effect of truncating terms in Equation  4 using the HMC based parameter inference. Truncating terms allows us to perform inference faster and can allow us to explore the tradeoff between dropping terms for speed and higher precision with more timesteps. In the right column of Figure  5, for a single field, we plot the mean predictions and the $15.9-84.1$ (1 sigma) percentiles computed using 200 samples with the approximate $-\log p_{\phi}(x_{0}|\theta)$ using Equation 4 as a function of $T_{\rm{MAX}}$ on the x axis for the same field. The true value of the cosmological parameter for this field is demarcated by the solid black lines. For this field and parameter, using more terms asymptotically removes the bias on $\Omega_{m}$ while increasing the bias on $\sigma_{8}$. However, using the first 20 timesteps only changes the mean prediction for $\Omega_{m}$ and $\sigma_{8}$ by $-0.26\%$ and $-1.27\%$ of the prediction using all 1000 timesteps respectively.
@@
We now turn our attention to the performance of our parameter inference approach relative to a power spectrum baseline. For subsequent HMC-based parameter inference in this section, we use Equation 4 with $T_{\rm{MAX}}=20$ to approximate the conditional negative log likelihood. Drawing 500 samples with $T_{\rm{MAX}}=20$ takes $\sim 32$ minutes for a single field.
To assess the constraining power of the power spectrum, we use the neural posterior estimation approach and train a normalizing flow to represent the posterior of $p_{\rm{NF}}(\theta|Pk\cdot\rm{Mean})$ using the Lampe [42] package. Since computing the overdensity power spectrum involves dividing by the mean of the field, we further concatenate the mean of the log field as an additional feature, since the prediction for $\Omega_{m}$ for a small box size of $25$ Mpc.$h^{-1}$ is very sensitive to the mean of the fields. For a single field, we compare the posteriors obtained by drawing 10000 samples from the power spectrum based estimator and 400 samples from the diffusion model+HMC in Figure  6. Other details about our implementation are in the Appendix  A.2.
@@
Figure 7: First row: An example field corresponding to $\Omega_{m},\sigma_{8}=0.33,0.96$ with its perturbed counterparts, along with the ratio of the perturbed linear field’s power spectrum to that of the original field $\frac{Pk_{Noisy}}{Pk_{Original}}$. Second, Third rows: Mean and $15.8-84.1$ percentile predictions using the Diffusion-HMC estimates (green) and the mean and sigma predictions obtained from the parameter inference network in [55], denoted by InfNet. The mean absolute percentage bias (of the true prediction) for each parameter is also indicated. The Diffusion-HMC constraints are more robust to perturbations to the input image.
IV.3 Robustness